# Predicting the Geothermal Gradient in Colombia: a Machine Learning Approach

### v1

@@ 5.3 Spatial Distribution of Predicted Geothermal Gradients
The spatial distribution of the predicted geothermal gradients is a key factor in assessing the utility of the model for geothermal exploration. Figure 11 illustrates the predicted geothermal gradients in the study area. The color gradient from blue to red indicates increasing values, with warmer colors representing higher geothermal potential.
Figure 11: Predicted geothermal gradient map showing the spatial distribution of geothermal potential across the region.
The map offers practical information on areas with significant geothermal activity, marking them as prime candidates for detailed exploration and potential development. The effectiveness of the predictive model lies in its ability to identify these spatial variations and key trends. This characteristic makes it a valuable asset in guiding exploration strategies towards the most promising geothermal energy locations.
5.3.1 Analysis of Predicted Gradients
The model’s output indicates a distinct sensitivity to geothermal indicators, revealing high-gradient values that are suggestive of underlying geothermal resources. These values are particularly pronounced in regions characterized by volcanic activity or where the crust is considerably thinner.
A standout feature of the model is its predictive accuracy in the Amazon basin, where data scarcity has previously been a barrier. The model’s gradient predictions for this area not only fill a critical data gap but also exhibit remarkable agreement with the findings from the Pimentel & Hamza (2012) study of Brazil’s portion of the basin. This concordance is a testament to the model’s robustness and its ability to extrapolate from known data to regions lacking in situ measurements.
In terms of new discoveries, the model highlights areas with potentially high geothermal gradients that are not documented in existing gradient maps. These predicted hotspots, particularly in the uncharted Amazon region, present opportunities for exploration and possible development of geothermal resources.
The model has proven to be a valuable asset in geothermal exploration, especially for the Amazon of Colombia, where it predicts high geothermal gradients in the absence of empirical data. Its alignment with the established geothermal gradients from Pimentel & Hamza’s work further validates its predictive capacity. As the model continues to be refined, its role in the identification and assessment of geothermal resources is expected to become critical, supporting a more informed and strategic approach to energy exploration in the region.
5.3.2 Implications for Geothermal Exploration
The spatial distribution map and the histogram of the predicted geothermal gradients provide critical guidance for geothermal exploration. These tools enable the strategic prioritization of regions for further investigation, allowing for a focused allocation of resources to areas with the highest geothermal potential.
The regions with the highest predicted gradients warrant immediate attention for exploration and data acquisition. These areas, flagged by the model, are likely candidates for sustainable geothermal development. In contrast, areas with lower predicted gradients can be investigated for baseline data collection, contributing to a more comprehensive understanding of the geothermal profile of the region.
Our predictive model can play a pivotal role in the early stages of exploration, for example, inform decision-making processes regarding site selection for geothermal power plants and drilling operations. By leveraging the model’s predictions, stakeholders can enhance their exploration strategies, targeting locations with a higher likelihood of geothermal activity, and thereby advancing the exploration for renewable energy sources. This optimized approach promises not only greater efficiency in resource utilization but also a higher success rate in developing geothermal energy projects.
@@ 6 Discussion
The evaluation of our predictive model for the estimation of geothermal gradients has been extensive, incorporating a variety of metrics and visual analyzes. Training set metrics, MAE, RMSE, and R-squared, along with the corresponding test set metrics provide a comprehensive view of the model performance.
@@
The model shows strong performance in the training set, with an MAE of 1.87, an RMSE of 2.50, and an R-squared value of 0.77, as indicated in Section 5.1. These values reflect a high degree of accuracy and a strong correlation between the predicted and actual geothermal gradients, suggesting that the model is capable of capturing the underlying patterns within the training data. However, when applied to the test set, the model’s performance drops, with an MAE of 2.61, an RMSE of 3.64, and an R-squared value of 0.51, as detailed in Section 5.1. The observed increase in error and the decrease in the $R^{2}$ value on the test set, as opposed to the training set, may suggest possible overfitting, particularly in the context of estimating higher values of geothermal gradients where the data are less abundant. Despite this, the model’s ability to generalize is demonstrated to be effective, as illustrated in Figure 4.
@@
Comparative analysis with related studies highlights the effectiveness of our model. For example, the study by Lösing & Ebbing (Op. Cit) in Antarctica achieved an nRMSE of 0.29 and an of $R^{2}$ 0.45, while Rezvanbehbahani et al. (OP. Cit) in Greenland recorded an nRMSE of 0.15 and an $R^{2}$ of 0.6. Our results, with an nRMSE of 0.12 and an $R^{2}$ of 0.52, not only reflect a lower relative error compared to these studies, but also show a balanced performance in terms of the determination coefficient. This comparison highlights our model’s accuracy and reliability across different geological settings.
The geospatial distribution of the predicted geothermal gradients, illustrated in Figure 11, highlights the spatial prediction capabilities of the model. The alignment of high-gradient regions in the actual vs. predicted maps (Figure 5) and the regions of significant discrepancy provide a nuanced understanding of the spatial accuracy of the model.
It is important to note that our model effectively predicted geothermal gradients in data-limited areas like south-western Colombia, including the Amazon and eastern Llanos, that were consistent with Pimentel & Hamza’s (2012) study, which observed higher gradients in the north-western Amazon basin. Training data shows an isolated low-gradient data point near the Amazon (Figure 5), however, that point has not influenced model training or prediction and can therefore be considered an outlier given the geological context of the south-west Colombia.
@@
Our feature importance analyses, comprising weight-based, gain-based, and mutual information methods, showcased the varying impact of different features. Elevation and basement proximity consistently emerged as the main predictors in multiple importance measures, as discussed in Sections 5.2. Moho depth is also one of the most important predictors, it´s relation with the lithospheric thickness suggests an agreement with the correlation of this tectonic feature with the thermal gradient defined by Kolawole & Evenick (2023). The geophysical properties, such as Bouguer or Free Air Anomaly, have been highlighted as significant by certain importance methods but appear less influential in others.
@@
The observed discrepancies between training and test performance may indicate the need of further model refinement or data enhancement. Strategies such as feature engineering, parameter optimization and integration of more specific data such as depth to basement or lithospere thickness could be employed to enhance the model’s predictive power and generalizability.
@@ 7 Conclusion
This study has successfully developed a machine learning model to estimate geothermal gradients across Colombia, marking a significant contribution in the field of geothermal energy exploration. By utilizing diverse geophysical and geological data, and gradient boosting regression; we have created a model that displays accurate gradient predictions across unexplored regions in Colombia.
Our model illustrates the potential of machine learning in enhancing the understanding of geothermal gradients distribution, specially within the context of Colombia, where there are vast regions whose geothermal energy has been underexplored (or not explored at all) due to the lack of extensive well data.
The incorporation of a wide range of geological and geophysical features was crucial to our approach, which allowed for a better understanding of the factors influencing geothermal gradients and heat distribution within the subsurface. The findings of this study mark a substantial step forward in geothermal exploration efforts, providing a promising tool for the identification of regions with significant geothermal potential. This advancement is expected to greatly assist in the prioritization of future geothermal exploration plans, making it a valuable asset for both researchers and industry professionals.
As we look to the future, the study encourages further research to refine and enhance the model’s accuracy. This includes the integration of local or regional data instead of global data, and the exploration of additional features that may influence geothermal gradients. Finally, we highlight the capability of Machine learning to contribute, through continued improvements, to the sustainable development of Colombia’s geothermal resources and other renewable energy sources.
@@ 8 Data Availability
The data supporting the findings of this study are openly available in the GitHub repository associated with this project. Interested parties can access the dataset, additional resources, and the code used for the analyses at the following URL:
This repository includes all relevant data, scripts, and additional materials required to replicate the study’s results or to further explore the geothermal gradient estimation model. By providing the data in a public domain, we encourage transparency, reproducibility, and collaborative efforts to enhance the research in geothermal gradient prediction using machine learning techniques.

### v7

@@ 6.3 Spatial Distribution of Predicted Geothermal Gradients
The spatial distribution of the predicted geothermal gradients is a key factor in evaluating the utility of the model for geothermal exploration. Figure 17 illustrates the geothermal gradients predicted in the study area. The color gradient from blue to red indicates increasing values, and warmer colors represent a higher geothermal potential.
Figure 17: Predicted geothermal gradient map showing the spatial distribution of geothermal gradient across the region.
The map offers practical information on areas with significant geothermal activity, marking them as prime candidates for detailed exploration and development; however, it’s important to understand the model limitations and, therefore, the fact that some high-gradient zones tend to be underrepresented. The effectiveness of the predictive model lies in its ability to identify these spatial variations and key trends, this characteristic making it a valuable asset in the guide of exploration strategies toward the most promising geothermal energy locations.
@@ 7 Discussion
The evaluation of our predictive model for the estimation of geothermal gradients has been extensive, incorporating a variety of metrics and visual analyses. Training set metrics, MAE, RMSE, and R-squared, along with the corresponding test set metrics provide a comprehensive view of the model performance. The model shows strong performance in the training set, with an MAE of 1.48, an RMSE of 2.14, and an R-squared value of 0.83, as indicated in Section 6.1. These values reflect a high degree of accuracy and a strong correlation between the predicted and actual geothermal gradients, suggesting that the model is capable of capturing the underlying patterns within the training data. However, when applied to the test set, the model’s performance drops, with an MAE of 2.68, an RMSE of 3.58, and an R-squared value of 0.55, as detailed in Section 6.1. The observed increase in error and the decrease in the $R^{2}$ value on the test set, as opposed to the training set, may suggest possible overfitting, particularly in the context of estimating higher values of geothermal gradients where the data are less abundant. Despite this, the model’s ability to generalize is demonstrated to be effective, as illustrated in Figure 9.
@@
Comparative analysis with related studies highlights the effectiveness of our model. For example, the study by Lösing & Ebbing (2021) in Antarctica achieved an nRMSE of 0.29 and an $R^{2}$ of 0.45, while Rezvanbehbahani et al. (2017) in Greenland recorded an nRMSE of 0.15 and an $R^{2}$ of 0.6. Our results, with an nRMSE of 0.12 and an $R^{2}$ of 0.55, reflect not only a lower relative error compared to these studies, but also a balanced performance in terms of the determination coefficient. This comparison highlights the accuracy and reliability of our model in different geological settings.
@@
Our feature importance analyses, comprising weight-based, gain-based, and mutual information methods, showcased the varying impact of different features. Elevation and proximity to the basement consistently emerged as the main predictors of multiple importance measures, as discussed in Sections 6.2. Moho depth is also one of the most important predictors, and its relation with the lithospheric thickness suggests an agreement with the correlation of this tectonic feature with the thermal gradient defined by Kolawole & Evenick (2023). The geophysical properties, such as Bouguer or Free Air Anomaly, have been highlighted as significant by certain importance methods but appear less influential in others.
@@
The observed discrepancies between training and test performance may indicate the need for further model refinement or data enhancement. Strategies such as feature engineering, parameter optimization, and integration of more specific data such as actual depth to basement or lithosphere thickness could be employed to enhance the model’s predictive power and generalizability. The geospatial distribution of the predicted geothermal gradients, illustrated in Figure 17, highlights the spatial prediction capabilities of the model. The alignment of high-gradient regions on actual versus predicted maps (Figure 11) and the regions of significant discrepancy provide a nuanced understanding of the spatial precision of the model.
Since the geothermal gradients used for training the model are apparent gradients and were averaged to build the original dataset, we must consider that limitation and treat both the preliminary geothermal gradients map (Figure 4a) and our predicted geothermal gradients map (Figure 17) as indicative tools rather than definitive measures. These maps provide a pragmatic framework for estimating geothermal potential, which is particularly useful to inform future exploration studies. This approach is especially beneficial in regions where well measurements are scarce or absent, which is the target of our study.
Despite the inherent limitations of the geothermal gradient data, which is depth-dependent and measured at varying depths, the model’s output demonstrates notable sensitivity to geothermal indicators, and reveals high-gradient trends that strongly suggest the presence of underlying geothermal resources. These values are particularly pronounced in regions characterized by volcanic activity or where the crust is considerably thinner. The alignment of high geothermal gradients with thermal manifestations such as volcanoes and hot springs, shown in Figure 3, and their association with fault lines detailed in Figure 1, underscores the influence of tectonic activities on these gradients. This may indicate geothermal systems dominated by convective heat transfer, especially in areas where faulting facilitates the ascent of heat from deeper sources within the earth.
Furthermore, the geothermal gradient data display responsiveness to more stable geological zones such as the Amazon and the Guiana shield. In these regions, the crust is relatively thinner and the Moho discontinuity is shallower (Figure 7), and the proximity of the basement to the surface enhances the efficacy of conductive heat transfer, signaling the presence of conduction-dominated geothermal systems. A standout feature of the model is its predictive accuracy in the Amazon, where data scarcity has previously been a barrier. The gradient predictions of the model for this area not only fill a critical data gap, but also exhibit remarkable agreement with the findings from the Pimentel & Hamza (2012) study of the Brazil’s portion of the Amazon (Figure 4d). This concordance is a testament to the model’s robustness and its ability to extrapolate from known data to regions lacking in situ measurements.
While the predicted map tends to underestimate high gradients, as seen in previous sections, we can infer that the geothermal gradient model is effectively capturing areas with geothermal potential, especially where geological structures such as faults are prevalent, and thermal manifestations are observed (refer to Figures 1, 2, and 3). This suggests that the geothermal model could serve as a valuable tool for targeting exploration efforts. Furthermore, the correlation between high-gradient areas and volcanic activities, as well as fault distributions and basement (igneous and metemorphic) rocks, could inform a more detailed geothermal map that includes both predicted gradients and direct geological measurements.
Comparison between our predicted map and the preliminary geothermal gradient map (Figure 4a) shows that, although the predicted map provides a competent representation of geothermal trends (as discussed previously), underestimating high gradients leads to significant differences in certain areas, such as the west of Eastern-Llanos and Caguán-Putumayo basins. This trend towards underestimation is likely attributed to an under-representation of higher gradient data in the training dataset. While we have attempted to mitigate this issue by assigning greater weights to these less frequent, high gradient instances (Figure 6), it is evident that acquiring and integrating additional high-gradient data is essential for enhancing the model’s accuracy in these critical areas. Further refinement in data collection and integration techniques will be crucial to improve the representation of high geothermal gradients by the predictive model.
Our predicted geothermal gradient map, when compared with the one produced by Gómez et al. (2019) for northwestern Colombia (Figure 4b), presents some discrepancies, particularly in very-low gradient values. These discrepancies could be attributed to an under-representation of low gradient values in our dataset, as well as to the general paucity of data in the region, as illustrated in Figure 5(a). The methodology followed by Gómez et al. (2019) might also contribute to the observed variances between their findings and the predictions of our model.
Conversely, when comparing our predicted geothermal gradient map with the one developed by Matiz-León (2023) for the Eastern-Llanos (Figure 4c), which utilized a geo-statistical methodology, there is a notable congruence in gradient distribution. It’s important to mention, however, that Matiz-León’s map indicates somewhat higher gradient values. This discrepancy may arise from a variety of factors, including the different methods applied (geo-statistical modeling and interpolation over a localized area versus our machine learning approach applied countrywide) as well as under-representation of high gradient data for training our model.
It is crucial, nevertheless, to acknowledge the variations observed between the maps of Matiz-León (2023) and Gómez et al. (2019) when compared to the preliminary geothermal gradient map by Alfaro et al. (2009). Such differences not only underscore the intrinsic challenges of estimating geothermal gradients in Colombia but also how different methodological frameworks can yield slightly different results. These discrepancies could be mitigated by enriching the dataset with more direct measurements, particularly in regions that are currently data-deficient. This would enable a more refined and accurate estimation of the country’s geothermal gradient landscape.
In terms of new discoveries, the model highlights areas with potentially high geothermal gradients that are not documented in existing gradient maps. These predicted hotspots, particularly in the uncharted Amazon region, present opportunities for exploration and possible development of geothermal resources. The model has proven to be a valuable asset in geothermal exploration, especially for the Amazon of Colombia, where it predicts high geothermal gradients in the absence of empirical data. Its alignment with the established geothermal gradients from Pimentel & Hamza’s (2010) work, and its similarity to other gradient maps, further validates its predictive capacity.
It is also important to recognize a fundamental limitation of our model, which is inherent to machine learning models that process vast amounts of data: the challenge of accurately predicting anomalously high geothermal gradients. For example, the over 65 °C/km gradients in specific locations at sedimentary basins, or the exceptionally high gradient at the Las Nereidas well in the Nevado del Ruiz Volcano in the Central Cordillera, which approaches 140°C/km as reported by Alfaro et al. (2009), are difficult to predict due to its rare occurrence in the dataset. Such outliers are typically treated as singular data points and may be excluded from the model’s training set to maintain generalizability. Specifically, in our analysis, we focused on a confidence range between the 1st and 99th percentiles, omitting these anomalous values to avoid skewing the model with disproportionately weighted data. Therefore, measuring anomalously high geothermal gradients accurately requires more constrained, direct, and localized methodologies, tailored to specific areas where these unusual conditions may occur.
Indeed, the application of data-driven methods like Machine Learning in predicting geothermal gradients is inherently reliant on the quantity and quality of available field data. This reliance emphasizes an ongoing opportunity for enhancement, as additional data can be used to continuously refine the model. Such improvements are expected to significantly enhance the model’s utility in the precise identification and assessment of geothermal resources. As the model continues to be refined, its role in the identification and assessment of geothermal resources is expected to become critical, supporting a more informed and strategic approach to energy exploration in the region.
The spatial distribution of the predicted geothermal gradients provide critical guidance for geothermal exploration. These tools enable the strategic prioritization of regions for further investigation, allowing for a focused allocation of resources to areas with the highest geothermal potential. The regions with the highest predicted gradients warrant immediate attention for exploration and data acquisition. These areas, flagged by the model, are likely candidates for sustainable geothermal development. In contrast, areas with lower predicted gradients can be investigated for baseline data collection, contributing to a more comprehensive understanding of the geothermal profile of the region.
Our predictive model can play a pivotal role in the early stages of exploration, for example, informing decision-making processes regarding site selection for geothermal power plants and drilling operations. By leveraging the model’s predictions, stakeholders can enhance their exploration strategies, targeting locations with a higher likelihood of geothermal activity, and thereby advancing the exploration for renewable energy sources. This optimized approach promises not only greater efficiency in resource utilization but also a higher success rate in developing geothermal energy projects. Finally, we hope that this study and, more importantly, the open training dataset we have provided, will motivate other researchers to explore additional machine learning and data-driven techniques, which without a doubt would lead to improved models.
@@ 8 Conclusion
This study successfully developed and validated a machine learning model using a Gradient-Boosted Regression Tree algorithm to predict geothermal gradients across Colombia. The effort marks a significant advancement in geothermal energy exploration, particularly within areas of Colombia that have previously lacked sufficient geophysical and geological data.
By employing a diverse array of geological and geophysical data, our model has demonstrated its capacity to provide accurate predictions of geothermal gradients. The accuracy of the model’s predictions has been rigorously validated against independent measurements, achieving an accuracy within 12% of these external data points. This level of accuracy substantiates the model’s effectiveness and supports its application in predicting geothermal gradient trends in regions that are underexplored or entirely unexplored due to the absence of direct measurements.
The application of machine learning techniques, specifically the Gradient-Boosted Regression Tree algorithm, has proved instrumental in synthesizing complex data into understandable and actionable insights. This approach has allowed us to generate a comprehensive geothermal gradient map of Colombia, which serves as a crucial tool in identifying regions with significant geothermal potential. The map serves as indicative for areas with elevated geothermal gradients, suggesting promising locations for future exploration and resource assessment.
The findings from this study can be particularly impactful for the geothermal energy sector in Colombia. The identified high-potential areas can guide future exploration efforts, ensuring that resources are allocated efficiently and effectively to locations most likely to yield substantial geothermal resources. This strategic focus is expected to significantly advance Colombia’s capabilities in geothermal energy production, supporting the country’s goals for sustainable energy development and contributing to global renewable energy efforts.
Looking forward, the study underscores the need for continued enhancements in machine learning methodologies to further refine the accuracy of geothermal gradient predictions. It is recommended that future research incorporate more localized and detailed geological data, which could provide deeper insights into subsurface conditions and improve model predictions. Additionally, exploring new geophysical features that might influence geothermal gradients could unveil further intricacies of geothermal systems.
Moreover, as the geothermal gradient data used in this study are derived from average measurements across varied depths, it is crucial to approach the interpretation of these data with an understanding of their inherent limitations. The integration of direct, high-resolution temperature and depth measurements could significantly enhance the accuracy of future predictions and provide a more detailed understanding of geothermal systems.
This research demonstrates the efficacy of Machine Learning tools for estimating geothermal gradients, particularly in the challenging context of Colombia. The accuracy and practicality of the model validate the value of data-driven approaches in enhancing exploration techniques with scarcity of direct measurements. Nevertheless, it is acknowledged that the algorithm has considerable scope for further improvement, particularly in the handling of the complexity of geothermal data and the enhancement of prediction accuracy across diverse geological settings. This work provides a robust foundation for the advancement of these techniques, with ongoing refinements anticipated to enhance the precision and effectiveness of these predictive models in geothermal resource assessment.
@@ 9 Data Availability
Data supporting the findings of this study are available in the GitHub repository associated with this project. Interested parties can access the dataset, additional resources, and the code used for the study at the following URL:
The complete data set of apparent geothermal gradient can be requested from the Servicio Geológico Colombiano (SGC) via their online assistance channel: https://www.sgc.gov.co/
10 Acknowledgements
We thank Servicio Geológico Colombiano (SGC) for generously collecting and providing the data used in this study. Manuel A. Florez received financial support from Vicerrectoría de Investigación y Extensión - Universidad Industrial de Santander, under Grant No. 3755.